OpenStudy (warpedkitten):

In triangle DEF, given the lengths of the sides, find the measure of the stated angle to the nearest degree. d = 5.4 e = 10.5 f = 10.8; m

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

use the law of cosines to find angle F f^2 = d^2 + e^2 - 2*d*e*cos(F) 10.8^2 = 5.4^2 + 10.5^2 - 2*5.4*10.5*cos(F) ... ... ... F = ??

OpenStudy (warpedkitten):

@jim_thompson5910 I keep using that, however I never get a straight answer. I use Mathway and it gives me 116.64=139.41-113.4cos(F) and then I'm not sure what to do with it.

jimthompson5910 (jim_thompson5910):

Let z = cos(F) for now. If we replace cos(F) with z, we get this equation 10.8^2 = 5.4^2 + 10.5^2 - 2*5.4*10.5*z are you able to solve for, or isolate, z?